In Country A, the population mean height for 3-year-old boys is 39 inches. Suppose a random sample of 15 3-year-old boys from Country B showed a sample mean of 38.4 inches with a standard deviation of 4 inches. The boys were independently sampled. Assume that heights are Normally distributed in the population. Complete parts a through c below.

Find the p-value.

p equals____? (Type an integer or decimal rounded to three decimal places as needed.
Answer #1

Solution:

For the given problem test statistic is:

t= -0.581

P-value= 2*p( t < -0.581)= 2*0.2855= 0.571

4, The provided sample mean is X = 38.4 and the sample standard deviation is s = and the sample size is n = 15. (1) Null and